OpenStudy (anonymous):

find an equation of the form (3) for the circle with the given equation: 2x^2+2y^2+2x+2y-1=0

OpenStudy (anonymous):

form (3)?

OpenStudy (anonymous):

the form 3 is just (x-h)^2+(y-k)^2=r^2

OpenStudy (anonymous):

oh ok first divide by 2 x^2 + y^2 + x + y = 0.5 (x + 0.5)^2 - 0.25 + (y + 0.5)^2 - 0.25 = 0.5 (x + 0.5)^2 + (y + 0.5)^2 = 1

OpenStudy (anonymous):

find the radius and the center point

OpenStudy (anonymous):

that last equation that @joeywhite wrote should give you the center and radius because it is what's called center-radius form.

OpenStudy (anonymous):

i know that the radius is 1^2 which is 1 but what is the center?

OpenStudy (anonymous):

compare the equation to the general form (x-h)^2+(y-k)^2=r^2 (x + 0.5)^2 + (y + 0.5)^2 = 1 the center is (h, k) and r is radius so for example the x coordinate of center is -h = + 0.5 h = -0.5